OpenStudy (fools101):

Help! Find the center and the radius of the circle. x2 + (y + 3)2 = 25 A) C(0, -3); r = 25 B) C(0, 3); r = 5 C) C(3, 0); r = 5 D) C(0, -3); r = 5

OpenStudy (anonymous):

i know first you factor

OpenStudy (turingtest):

\[(x-h)^2+(y-k)^2=r^2\]is a circle radius r with center at (h,k) identify h and k and r, then you have your answer

OpenStudy (turingtest):

no factoring necessary ;)

OpenStudy (anonymous):

huh really because you have a 2(y+3) dont you need to distrubute that first

OpenStudy (lalaly):

\[x^2+(y+3)^2=25\]

OpenStudy (turingtest):

lol, yes it would seem that way, but I know they mean \[x^2 + (y + 3)^2 = 25\] otherwise it would not be a circle. people get lazy and just copy and paste questions, so it looks like a coefficient.

OpenStudy (fools101):

it B right?!

OpenStudy (anonymous):

oh its an exponent

OpenStudy (anonymous):

shoot

OpenStudy (lalaly):

its D

OpenStudy (fools101):

why is it D -3? i thought it would be B isn't the 3 positive?

OpenStudy (turingtest):

careful about the minus sign!\[x^2+(y+3)^2=25=x^2+(y-(-3))^2=25\]so k=-3

OpenStudy (lalaly):

look at what turingtest wrote at the begining 25 = r^2 so r=5 and the general form of the circle is\[(x-h)^2+(y-k)^2=r^2\] the center is (h,k)

OpenStudy (fools101):

AWWW okay i see i see!

OpenStudy (fools101):

thank u!!